Collins limited is considering investing K40 million in equipment which will generate a net cash flow of K16 Million per year four years. The company is able to depriciate the equipment at the rate of of 20% per year on a straight line for tax purposes. The market value of the equipment at the end of the four years is expected to be K15 million. The difference between the market value and the equipments tax value(cost less depreciation to date of sale) is termed as recoupment which in this case is subject to tax.

The corporation tax is 28%. the companys cost of capital is 14%.

Assume depreciation is allowable as deduction before computing tax and that the tax as an outflow from the cash at 28%

a)What is the projects net prsent value

b) What is the projects IRR

C) What is the projects payack

d) Is the project viable? justify your deisio
Answer #1

Depreciation of the equipment = 20% of K40 million = K8 million per year. Hence, the equipment's depreciable life is 5 years.

Inflow per year= K16 million

Tax = 0.28 (K16 million -depreciation) = 0.28 (K16 million - K8 million) = K2,240,000

Cash flow per year= K16,000,000 - tax + add back non-cash items (depreciation) = K21,760,000

At the end of the fourth year, the equipment's book value is K40 million - 4* (K8 million) = K8 million

The market value of the equipment at the end of the 4th year is K15 million. The excess over book value = K15 million -K8 million = K7 million.

Tax on sale of equipment is K7 million * 0.28 = K1.96 million.

Now, we need to bring back all these values back in present terms by discounting it by 14%

NPV = (K21,760,000/1.14) + {K21,760,000/(1.142)} + {K21,760,000/(1.143)} + {K21,760,000/(1.144)} +  {K15,000,000/(1.144​​​)} - {K1,960,000/(1.144​​​​​​​)} - K40,000,000 = K31,123,106.56

IRR = 45.60% (calculate using financial calculator)

Payback period = cost of equipment/ yearly cash flow = K40,000,000 / K21,760,000 = 1.84 years or just over 1 year and 10 months

The project is viable, the NPV of the project is quite high, and so is the IRR. Meanwhile, the payback period is within 2 years, which is more than acceptable, considering that the project is being planned for the next 4 years